A electric beam is incident on a potential barrier at height 0.03ev and of infinite with find the reflection coefficient of the barrier if the energy of the incident electric beam be (a)0.04ev(b)0.025ev(c)0.03ev.


1
Expert's answer
2021-03-24T19:58:18-0400

GIVEN:

"V_o=0.03eV
\\a) E_1=0.04\\b)E_2=0.025\\c)E_3=0.03"

first we find the transmission coefficient (T) for the step like potential:

"\boxed{T={4\sqrt E \sqrt{E-V_o}\over (\sqrt E + \sqrt{E-V_o})^2}}" ............1


then reflection coefficient (R).

"\boxed{R=1-T}" ...............................2


now using 1 and 2 we calculate "T_1,T_2,T_3" AND "R_1,R_2,R_3"

a)

"\boxed{T_1={4\sqrt 0.04 \sqrt{0.04-0.03}\over (\sqrt0.04 + \sqrt{0.04-0.03})^2}=0.889}
\\"


"\boxed{R_1=1-T=0.111}"

b)

"\boxed{T_2={4\sqrt 0.025 \sqrt{0.025-0.03}\over (\sqrt0.025 + \sqrt{0.025-0.03})^2}=error}
\\" no value exists because ,E<V and squre root is not defined for negative value.

therefore, value of "R_2=" not defined


c)

"\boxed{T_3={4\sqrt 0.03 \sqrt{0.03-0.03}\over (\sqrt0.03 + \sqrt{0.03-0.03})^2}=0}"


therefore

"\boxed{R_3=1-0=1}"